Determine the IRR of the following:

A) An initial investment of $10,750 resulting in a free cash flow of $3,500 at the end of year 1, $2,000 at the end of year 2, and $7,000 at the end of year 3.

B) An initial investment of $5,000 resulting in a SINGLE free cash flow of $38,077 after 10 years.
Answer #1

Let irr be x%
At irr,present value of inflows=present value of outflows.

a.

10,750=3,500/1.0x+2,000/1.0x^2+7,000/1.0x^3

Hence x=irr=6.92%(Approx).

b.5,000=38,077/1.0x^10

1.0x^10=(38,077/5000)

1.0x =(38,077/5000)^(1/10)

x=1.2251-1

=22.51%(Approx).